2) Determine whether the function rule models discrete or continuous data.
A produce stand sells roasted peanuts for $2.99 per pound. The function C(p) = 2.99p relates the total cost of the peanuts to the number of pounds purchased p.
A) Discrete
B) Continous
6) Choose the correct function rule for values in the table.
x f(x)
2 –8
3 –12
4 –16
5 –20
A) f(x) = -4x
B) f(x) =4x
C) f(x) =x-4
D) f(x)=x+4)
7) Choose the correct function rule for values in the table.x f(x)
X f(x)
3 7
4 8
5 9
6 10
A) f(x) =x-4
B) f(x)=4x
C) f(x) =x+4
D) f(x)= -4-x
8) The length of a field in yards is a function f(n) of the length n in feet. Choose the function rule that models this situation.
A) f(n)=1/3n
B) f(n)=3n
C) f(n)=12n
D) f(n)=n/12
9)Choose the function rule that gives the total cost c(p) of p pounds of sugar if each pound costs $0.59.
A) c(p)=59p
B)c(p)=p/0.59
C)c(p)=p+0.59c
D)c(P)=0.59
10) Choose the equation of the direct variation that includes the point (9, 12).
A) Y= -1/1/3x
B) y=-1/12x
C) y=1/1/3x
D) y=-3/4x
pleae help me...it's my QUIZ :((

Answer by jim_thompson5910(35256)

I'll do the first few to get you going.

# 2
The key words are "per pound", which means that this is a discrete variable.

So the answer is choice A
=======================================================
# 6
As x increases by 1, f(x) increases by 4. So the slope is 4/1 = 4. So the equation is f(x) = 4x (since this is the only choice with a slope of 4)
=======================================================
# 7

As x increases by 1, f(x) increases by 1. So the slope is 1/1 = 1. So the equation is f(x) = x + 4 (since this is the only choice with a slope of 1 and plugging in x = 3 gives you 7)
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------